# Optical absorption in tilted geometries as an indirect measure of longitudinal plasma waves in layered cuprates

### v1

@@ IV Fresnel equations at normal incidence on a tilted-grown sample
In Fig. 3(b) we show the real part of Eq. (30) as a function of the external polarization angle and we compare the peak emerging in the measured conductivity with Eq. (32). Indeed, the approximated expression (31) provides an excellent description of the experimental data in Ref.s [34, 33, 35, 36], and the frequency Eq. (32) establishes a link between the peak of the experimental conductivity and the plasma frequencies $\omega_{ab}$ and $\omega_{c}$, which can then be extracted as fitting parameters given the angles $\eta$ and $\phi$. In Fig. 3(c) we fit experimental data from Ref. [36] to provide an estimate of the in-plane and out-of-plane plasma frequencies of the overdoped $\text{La}_{1.87}\text{Ce}_{0.13}\text{CuO}_{4}$ ($T_{c}=21$ K) at $5$ K. So far, it was only empirically observed in Ref. [33] that the data could be well fitted by using an effective conductivity $\sigma_{t}(\omega,\eta_{\text{eff}})$ having the same functional form of Eq. (4), but with an effective tilt angle $\eta_{\text{eff}}=\eta\sin\phi$. This result actually follows from Eq. (31) in the case of small angle $\eta$ between the momentum and the $c$-axis of the crystal, which is indeed the configuration of Ref. [33]. In this case, the frequency of the peak in Eq. (32) can be approximated as
@@ Acknowledgements
A propagating solution is allowed whenever the determinant of this system is zero. For the experimental configuration of Fig. 4(a) in which the frame $(x,y,z)$ corresponds to $(t,b,l)$, as in Section 3.3, the incoming momentum is along the $z$ (or $l$) direction. In this geometry the dielectric tensor is defined as (see Eq.s (2) and (17))
@@
as in Eq. (23) in the main text. Conversely, for the experimental configuration of Fig. 4(b) in which the frame $(x,y,z)$ corresponds to the crystallographic frame $(a,b,c)$, the dielectric tensor is diagonal $\hat{\varepsilon}_{\alpha\beta}=\varepsilon_{\alpha}\delta_{\alpha\beta}$ (see Eq.s (1) and (12)). As the wave-vector belongs to the $xz$-plane, again in this case the equation for $\text{E}_{y}$ is decoupled from the other two components. One can then in both cases solve the system separately for the transmission of the $y-$ and for the mixed $xz-$ polarized components of $\mathbf{E}$. In the following we will refer to the former component as the Transverse Electric (TE) field and to the latter as the Transverse Magnetic (TM) field, as one would commonly do in the oblique incidence configuration in which the plane of incidence is $xz$. Notice that the transmitted wave $\textbf{E}^{\text{T}}$ is generically polarized along a direction $\phi^{\prime}$ that differ from the polarization $\phi$ of the incident wave $\textbf{E}^{in}$, although the experiment is still set so to measure the outgoing $\phi$-polarized wave $\textbf{E}^{\text{T}^{\prime}}$. Assuming that the TE and TM transmission coefficients are known, such that $\text{E}^{\text{T}}_{\text{TE,TM}}=\text{T}_{\text{TE,TM}}\text{E}^{in}_{\text{TE,TM}}$ (see Fig. 5 for the notation), one finds the measured transmitted field polarized along $\phi$ as
@@
Even though Eq.s (47) and (48) still depend on a combination of $\varepsilon_{ab}$ and $\varepsilon_{c}$, these structures do not lead to the pole observed in the transverse dielectric function $\varepsilon_{t}$, as opposed to Eq. (41) obtained in the first configuration. In the end, by explicit numerical computation with realistic parameter values for cuprates of the transmissivity $\text{T}_{\text{TM}}$ and $\text{T}_{\text{TE}}$ in Eq. (45), with the corresponding definitions of $\eta$, we verified that the corresponding conductivities, expressed as in Eq.  (43), are featureless, and no finite-frequency peaks are observed.

### v2

@@ IV Fresnel equations at normal incidence on a tilted-grown sample
In Fig. 3(b) we show the real part of Eq. (30) as a function of the external polarization angle and we compare the peak emerging in the measured conductivity with Eq. (32). Indeed, the approximated expression (31) provides an excellent description of the experimental data in Refs. [34, 33, 35, 36], and the frequency Eq. (32) establishes a link between the peak of the experimental conductivity and the plasma frequencies $\omega_{ab}$ and $\omega_{c}$, which can then be extracted as fitting parameters given the angles $\eta$ and $\phi$. In Fig. 3(c) we fit experimental data from Ref. [36] to provide an estimate of the in-plane and out-of-plane plasma frequencies of the overdoped $\text{La}_{1.87}\text{Ce}_{0.13}\text{CuO}_{4}$ ($T_{c}=21$ K) at $5$ K. We here clarify that the experimental fit is not meant to draw any conclusion on the symmetry of the superconducting order parameter, that is still debated in the context of electron-doped cuprates [36, 52]. Our starting point Eq. (9) can be derived from a microscopic model which admits a modulation of the superconducting gap [19]. At the level of Eq. (9), the gap symmetry enters in the temperature dependence of the superfluid stiffness $D_{ab}$ and $D_{c}$, that is controlled by quasiparticle excitations, while it barely affects the charge compressibility $\kappa_{0}$. In addition, the gap symmetry can affect the quasiparticle damping of the plasmon, controlling the phenomenological broadening $\gamma$, even though other mechanism can determine its value independently of the gap symmetry. However, once the proper gap symmetry has been embedded in the plasma frequencies $\omega_{ab}$ and $\omega_{c}$, the structure of the modes Eq. (18) is general.
So far, it was only empirically observed in Ref. [33] that the data could be well fitted by using an effective conductivity $\sigma_{t}(\omega,\eta_{\text{eff}})$ having the same functional form of Eq. (4), but with an effective tilt angle $\eta_{\text{eff}}=\eta\sin\phi$. This result actually follows from Eq. (31) in the case of small angle $\eta$ between the momentum and the $c$-axis of the crystal, which is indeed the configuration of Ref. [33]. In this case, the frequency of the peak in Eq. (32) can be approximated as
@@ Acknowledgements
A propagating solution is allowed whenever the determinant of this system is zero. For the experimental configuration of Fig. 4(a) in which the frame $(x,y,z)$ corresponds to $(t,b,l)$, as in Section 3.3, the incoming momentum is along the $z$ (or $l$) direction. In this geometry the dielectric tensor is defined as (see Eqs. (2) and (17))
@@
as in Eq. (23) in the main text. Conversely, for the experimental configuration of Fig. 4(b) in which the frame $(x,y,z)$ corresponds to the crystallographic frame $(a,b,c)$, the dielectric tensor is diagonal $\hat{\varepsilon}_{\alpha\beta}=\varepsilon_{\alpha}\delta_{\alpha\beta}$ (see Eqs. (1) and (12)). As the wave-vector belongs to the $xz$-plane, again in this case the equation for $\text{E}_{y}$ is decoupled from the other two components. One can then in both cases solve the system separately for the transmission of the $y-$ and for the mixed $xz-$ polarized components of $\mathbf{E}$. In the following we will refer to the former component as the Transverse Electric (TE) field and to the latter as the Transverse Magnetic (TM) field, as one would commonly do in the oblique incidence configuration in which the plane of incidence is $xz$. Notice that the transmitted wave $\textbf{E}^{\text{T}}$ is generically polarized along a direction $\phi^{\prime}$ that differ from the polarization $\phi$ of the incident wave $\textbf{E}^{in}$, although the experiment is still set so to measure the outgoing $\phi$-polarized wave $\textbf{E}^{\text{T}^{\prime}}$. Assuming that the TE and TM transmission coefficients are known, such that $\text{E}^{\text{T}}_{\text{TE,TM}}=\text{T}_{\text{TE,TM}}\text{E}^{in}_{\text{TE,TM}}$ (see Fig. 5 for the notation), one finds the measured transmitted field polarized along $\phi$ as
@@
Even though Eqs. (47) and (48) still depend on a combination of $\varepsilon_{ab}$ and $\varepsilon_{c}$, these structures do not lead to the pole observed in the transverse dielectric function $\varepsilon_{t}$, as opposed to Eq. (41) obtained in the first configuration. In the end, by explicit numerical computation with realistic parameter values for cuprates of the transmissivity $\text{T}_{\text{TM}}$ and $\text{T}_{\text{TE}}$ in Eq. (45), with the corresponding definitions of $\eta$, we verified that the corresponding conductivities, expressed as in Eq.  (43), are featureless, and no finite-frequency peaks are observed.